OpenStudy (anonymous):

Choose the value of the y determinant (Dy) in the following system. 3x – y = 7 2x + 5y = 4 17 39 28 –2

OpenStudy (anonymous):

The y determinant involves replacing the y coefficients with the constant terms. Dy = | 3 7 | = (3)(4) - (7)(2) = 12 - 14 = -2 | 2 4 |

OpenStudy (anonymous):

does that make sense? It's like the regular determinant, but in the 2nd column, you put in 7 and 4 in place of the y-coefficients.

OpenStudy (anonymous):

so it's -2? thnx, bro! :)

OpenStudy (anonymous):

glad to help :)